Pistenbully problems – University of Copenhagen

Forward this page to a friend Resize Print Bookmark and Share

EastGrip - The East Greenland Ice-core Project > Field Diaries Folder > UK diaries 2017 > Pistenbully problems

15 May 2017

Monday 15th of May 2017

From the drillers workshop. Bruno is telling Steff and Dennis how he has repaired the power plug for the drill.

No expedition without problems.

We tracked down the problem we have had with one of our Pistenbullys. It has a leak in the hydraulic brake that propagates in to the drive wheel. We have been in contact with the company, and we need a specialist up to do repairs.

Also today, the drillers had trouble with a specialized electrical plug on the drill. The plug connects the power and digital signal from the cable with the electronics inside the pressure compartment of the drill. All electronic of the drill has to be inside the pressure compartment as pressure of the liquid filled borehole will reach more than 200 bar when we drill down to 2.5 km and electronical components do not like this pressure. A little dirt between two connectors of the plug caused a short, and the plug had to be replaced.

I the mean time the drillers worked on preparing the drilling fluid handling system. They have invented a new method to reclaim the drilling fluid that hopefully will reduce our consumption of fluid. For people working on the surface, this afternoon became very nice as the wind calmed down a bit.

What we did today:

  1. Grooming apron and taxiways.
  2. Work in the drilling trench.
  3. Markers on the NE lead-in line have been reset. This completes resetting markers for this Summer.
  4. Test of radar.
  5. Making tags and labels for ice cores.
  6. Survey from 2016 reference markers to top of borehole lid in drill trench.
  7. Excavating the PROMICE weather station.
  8. Making ice core trough holders for shelf system in core buffer.
  9. German snow sampling.
  10. Japanese aerosol sampling and snow experiments in progress.
  11. Water vapour sampling, methane sampling, boundary layer and meteorology station working.

Ad.1: All skiway markers have been reset and checked. The lead-ins are now straight again. The skiway is even and hard and ready to receive a plane. We have fuel to give and we can take fuel. We have 8 ATO and igniters in camp.

Ad.2: We made a camera inspection of the borehole and prepared for a last reaming run. A short developed in the top electrical socket of the electronics section. Bruno replaced the socket and Jens Christian modified the firmware and software. Others worked on preparing the drilling fluid handling station.

Ad.6: The top of the lid of the borehole is 13.72 m below the May 2016 snow surface datum. This confirm measurements from last year and the top of the ice core drilled in the main hole is the intersection between bag 25 and 26 (13.75 m) . First bag of EGRIP main core is bag 26, complete.

Weather today: Sunshine all day. Temp. -18 °C to -29 °C. Wind: 10 kt from SW. Visibility: To horizon.

FL, J.P. Steffensen

← Previous entry   Next entry →